OpenStudy (anonymous):

find derivative. g(x)=(1+4x)^5(3+x-x^2)^8

OpenStudy (anonymous):

i need help with the steps. i know i have to do the chain rule twice but im having trouble simplifying

OpenStudy (anonymous):

You got to do the chain rule and the product rule. Let me work it out and i will post the steps

OpenStudy (anonymous):

i got this far. \[\left( 1+4x \right)^58\left( 3+x-x^2 \right)^7(1-2x)+\left( 3+x-x^2 \right)^85\left( 1+4x \right)^4(4)\]

OpenStudy (anonymous):

OH! common factor of [4(1+4x)^4(3+x-x^2)^7\] so [4(1+4x)^4(3+x-x^2)^7\] [(1+4x)^12(1-2x)+(3+x-x^2)^15\]

OpenStudy (anonymous):

you can factor out a \[(1+4x)^4(3 + x -x^3)^7\]leaving you with \[(1+4x)^4(3 + x -x^3) *[8(1+4x) (1-2x) + 20(3 +x -x^3)\] then i believe all you have to do is simplify the expression that you where left after you factored out what was common and you should have your answer

OpenStudy (anonymous):

common factor \[4\left( 1+4x \right)^4\left( 3+x-x^2 \right)^7\]

OpenStudy (anonymous):

ahh i forgot about the 4.. sorry lol

OpenStudy (anonymous):

ah yes, i can simplify the rest. i got 17+9x-21x^2

OpenStudy (anonymous):

yeah that should be your answer \[4(1+4x)^4(3 + x -x^3)(17 +9x -21x^2)\]

OpenStudy (anonymous):

NICE.!!

OpenStudy (anonymous):

i forgot the the exponent 7 to the (3 + x - x^3)

OpenStudy (anonymous):

oh pellet look at my next one

OpenStudy (anonymous):

haha i cant say s.h.i.t

OpenStudy (anonymous):

it changes s.h.i.t to pellet??? lol

OpenStudy (anonymous):

try it

OpenStudy (anonymous):

this pellet is crazy

OpenStudy (anonymous):

haha wow